Skip to content
dronacharya.ai
← JEE Advanced 2016 · Paper 1
MathematicsSingle correct+3 / -1Question 40 of 54

JEE Advanced 2016 · Paper 1, Mathematics question 40

A computer producing factory has only two plants T1T_1 and T2T_2. Plant T1T_1 produces 20% and plant T2T_2 produces 80% of the total computers produced. 7% of computers produced in the factory turn out to be defective. It is known that

PP(computer turns out to be defective given that it is produced in plant T1T_1) =10P=10P(computer turns out to be defective given that it is produced in plant T2T_2),

where P(E)P(E) denotes the probability of an event EE. A computer produced in the factory is randomly selected and it does not turn out to be defective. Then the probability that it is produced in plant T2T_2 is

Choose one option
Answer and worked solution

Official answer: C

Defect rates. Let P(D∣T2)=pP(D|T_2)=p, so P(D∣T1)=10pP(D|T_1)=10p. Then
0.07=0.2(10p)+0.8p=2.8p  ⟹  p=0.025.0.07=0.2(10p)+0.8p=2.8p\implies p=0.025.

Not defective.

  • P(Dˉ∣T1)=0.75P(\bar D|T_1)=0.75 and P(Dˉ∣T2)=0.975P(\bar D|T_2)=0.975.
  • P(Dˉ)=0.2(0.75)+0.8(0.975)=0.15+0.78=0.93P(\bar D)=0.2(0.75)+0.8(0.975)=0.15+0.78=0.93.

Bayes.
P(T2∣Dˉ)=0.780.93=7893.P(T_2|\bar D)=\frac{0.78}{0.93}=\frac{78}{93}.

(C)